OpenStudy (abbycross167):

Write the slope-intercept (y = mx + b) form of the equation from the information given. The line goes through the points (-1, -7) and (-8, -2) Slope = y2 - y1 =-1 - (-2) x2 - x1 -7 - (-8) = 2/1 @AbdullahM this is all I could figure out

OpenStudy (anonymous):

hows it going need some help?

OpenStudy (abbycross167):

yes sir/ma'am I don't know how to finish this problem

OpenStudy (anonymous):

Im A SIR!!! >_< lol dont worry, lets work out the problem... so i see you've tried to find the slope.. :)

OpenStudy (abbycross167):

yes sir... but I don't know how to get the equation

OpenStudy (anonymous):

thats alright the only thing is... you got the slope wrong..

OpenStudy (abbycross167):

oh... what would it have been? 2/2?

OpenStudy (anonymous):

no so if we plug in the numbers we get, -2-(-7)/-8-(-1) correct?

OpenStudy (anonymous):

@abbycross167

OpenStudy (abbycross167):

Yes sir

OpenStudy (abbycross167):

would it be 5/-7?

OpenStudy (anonymous):

correct :) so then we now have the slope

OpenStudy (abbycross167):

really? Awesome!

OpenStudy (anonymous):

yes so now we find the y-intercept....

OpenStudy (abbycross167):

How would I do that?

OpenStudy (anonymous):

we plug in our point (x,y), and solve for b. y=mx+b y=-5/7x+b

OpenStudy (abbycross167):

So y=-5/7x+b would be the equation

OpenStudy (anonymous):

yes, so we plug in one point, and solve for b. give it a try :)

OpenStudy (abbycross167):

ok wait... I'm confused (I'm sorry I'm not good at math) would y=-5/7x+b be the equation or I have to solve for b in order to get the equation?

OpenStudy (anonymous):

its fine im not rushing, math isnt easy for everyone but you need b to get the equation your looking for,, sorry that was my bad haha

OpenStudy (abdullahm):

Hi :)

OpenStudy (abdullahm):

@abbycross167 y2 - y1 would be -7 - (-2) = -7 + 2 = -5 x2 - x1 = -1 - (-8) = -1 + 8 = 7 So the slope is -5/7

OpenStudy (abbycross167):

Howdy... I really don't know what b would be

OpenStudy (abdullahm):

Now we plug in a point y = -5/7x + b -7 = -5/7 ( -1) + b Solve for b :)

OpenStudy (anonymous):

hows it going?@AbdullahM its fine, so of we plug in our point, we get -7=-5/7(-1)+b or b=-7-5/7(-1)

OpenStudy (anonymous):

ahh you beat me to it lol

OpenStudy (abbycross167):

would b be 5/7?

OpenStudy (anonymous):

haha i would actually use a calculator for this, not gonna lie, but i got, -54/7

OpenStudy (abbycross167):

@AbdullahM what did you get for b?

OpenStudy (abbycross167):

Would y=-5/7x+(-54/7) be the equation?

OpenStudy (anonymous):

yes that would be your final equation

OpenStudy (abbycross167):

oh ok, thank you kindly... would you mind helping me with one more?

OpenStudy (anonymous):

no problem.. sure ill give it a shot :)

OpenStudy (abbycross167):

Ok, would you like me to do a new post?

OpenStudy (anonymous):

nah, here's fine

OpenStudy (abdullahm):

Good job! That's correct :)